OpenStudy (anonymous):

Please explain me how to graph this : y= 3^x

OpenStudy (anonymous):

You have to look at the behavior of this graph at x approaching +-infinity and x = 0. If you know calculus, you can take the derivatives into account and realize that this function is always increasing and concave up. at x going to -infinity, the graph approaches 0 but is still positive, so the x-axis is a horizontal asymptote. At x=0, we have y=1. At x approaching +infinity, the graph goes to infinity.

OpenStudy (homeworksucks):

or just get a table of values and plot points but what @tcarroll010 said.

OpenStudy (anonymous):

Here's a better graph than the one I drew.

OpenStudy (anonymous):

Notic how 3^0 = 1 and 3^1 = 3. You would see that 3^2 = 9

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

@burhan101 ?

OpenStudy (anonymous):

how would you use a table of values ?

OpenStudy (anonymous):

i am not allowed to use any calculus to solve, it's an advanced function's course and my teacher does not allow :(

OpenStudy (anonymous):

You can generate a table of values by selecting various periodic values of x. Just plug in for x = -1000, -100, -10, -1, 0, 1/4, 1/2, 1, 2, 3 That should be enough values to get that curve and then just make a smooth curve for the graph.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

let me try this on paper, ill tell you if i get it

OpenStudy (anonymous):

You're welcome! Good luck in all your studies and thx for the recognition! @burhan101

OpenStudy (anonymous):

Here's a graph with some points

OpenStudy (anonymous):

looking at the base of the function, can i assume there will be a point at 1,3 ?

OpenStudy (anonymous):

yes, because 3^1 = 3 It will be shown as 1 of the points I drew.

OpenStudy (anonymous):

ohh okay ! so if the function was let's say y=5^x there would be a point at 5,1